# Optimal Algorithm for the Planar Two-Center Problem

### v3

@@ 2 Preliminaries
For any $r>0$, an $r$-solution for a planar two-center instance $S$ is a set of two radius-$r$ disks in the plane whose union covers $S$; a solution for $S$ refers to an $r$-solution with some $r\geq 0$. We denote by $\mathsf{opt}(S)$ the minimum $r$ such that $S$ has a $r$-solution. An $r$-solution $\{D,D^{\prime}\}$ for $S$ is tight if $\mathsf{dist}(D,D^{\prime})$ is minimized (among all $r$-solutions). An $r$-solution $\{D,D^{\prime}\}$ for $S$ is $p$-anchored for a point $p\in\mathbb{R}^{2}$ if $p\in D\cap D^{\prime}$. The anchored two-center problem takes as input a set $S$ of points in the plane and a point $p\in\mathbb{R}^{2}$, and asks for a $p$-anchored $r$-solution for $S$ with the minimum $r$. As discussed in Section 1, Choi and Ahn [8] presented an $O(n\log n)$-time algorithm for the anchored problem.
@@ 3 $\bm{r}$-Coverage
First of all, by the definition of $\alpha_{r}(Q)$, such a disk $D$ must exist for $p\in\partial\alpha_{r}(Q)$. In this proof, we let $q$ be the antipodal point $p$ on $\partial D$, and show that $q$ lies on $\partial\mathcal{CR}_{r}(Q)$.
@@
Let $e$ be an arc of $\alpha_{r}(Q)$ with two vertices $u$ and $v$. By slightly abusing the notation, define $\tau(e)=\cup_{p\in e}\tau(p)$. Let $D$ be the unique radius-$r$ disk whose boundary containing $e$. Note that $D$ contains $\alpha_{r}(Q)$ and thus contains $Q$ [13, 18, 29]. This property observes that $\tau(e)$ consists of the antipodal points of all points of $e$ on $\partial D$. We call $\tau(e)$ the antipodal arc of $e$. Since we have assumed that the distance of any two points of $Q$ is less than $2r$, $\tau(e)$ does not intersect $e$ [13, 18, 29]. For the reference purpose later, we have the following corollary, following directly from Lemma 6.
@@
We proceed to show that such a disk $D$ is unique. Assume there is another radius-$r$ disk $D^{\prime}$ with $Q\subseteq D^{\prime}$ and $q\in\partial D^{\prime}$. Then the antipodal point $p$ of $q$ on $\partial D$ lies outside of $D^{\prime}$. On the other hand, since $p$ lies on $\partial\alpha_{r}(Q)$ (which has been proved above), $p$ must be contained in $D^{\prime}$ because $D^{\prime}$ contains $Q$ (and thus contains $\alpha_{r}(Q)$). This incurs contradiction. ∎
@@ 4 Algorithm of Theorem 1
Clearly, $|A|=O(1)$. Roughly speaking, we uniformly sample $O(1)$ points in the radius-$2$ disk $x^{2}+y^{2}\leq{4}$ and put them in $A$; we also add $o$ to $A$.
@@ 5.1 Geometric lemmas
Let $\{D,D^{\prime}\}$ be a tight $r$-solution for $S$, and $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting the two disk centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. Then, there exist two points $u,v\in(S\cap\partial D)\backslash D^{\prime}$ such that $\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$; see Figure 4. Similarly, there exist two points $u^{\prime},v^{\prime}\in(S\cap\partial D^{\prime})\backslash D$ such that $\overline{u^{\prime}v^{\prime}}\cap(\ell\backslash\sigma)\neq\emptyset$.
@@
Let $\{D,D^{\prime}\}$ be a set of two radius-$r$ disks, $u,v\in\partial D\setminus D^{\prime}$ be two points whose antipodal points on $\partial D$ are not in $D^{\prime}$, and $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. If $\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$, then $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$.
@@
Let $\tau^{\prime}$ be the antipodal arc of $\tau$ on $\partial D$, i.e., $\tau^{\prime}$ consists of the antipodal points on $\partial D$ of all points of $\tau$. Note that $u$ and $v$ are two endpoints of $\tau^{\prime}$. Since $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$, $\tau^{\prime}$ is in a half circle of the circle containing $\tau^{\prime}$, and thus $\tau^{\prime}$ must be an arc of the circular hull $\alpha_{r}(\{u,v\})$ of $\{u,v\}$. Applying Corollary 7 with $Q=\{u,v\}$, we obtain that $\tau\subseteq\partial\mathcal{CR}_{r}(\{u,v\})$.
@@ 5.2 Proving the correctness
  • If $r<1$, since $\mathsf{dist}(D,D^{\prime})\leq(2-\frac{5}{c})\cdot r$, there exists a disk $D^{\prime\prime}\subseteq D\cap D^{\prime}$ of radius $1/c$, which in turn contains an axis-parallel square $\Box$ of side-length $1/c$. Note that one of $D$ and $D^{\prime}$ must intersect the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$, because $S\subseteq D\cap D^{\prime}$. As $r<1$, we know that $\Box\subseteq\{(x,y):x^{2}+y^{2}\leq 4\}$. By the same argument as above, $\Box$ contains some point $p\in\{(x,y)\in\mathbb{Z}_{/c}^{2}:x^{2}+y^{2}\leq 4\}\subseteq A$, and thus $\{D,D^{\prime}\}$ is $p$-anchored.

### v4

@@ 2 Preliminaries
For any $r>0$, an $r$-solution for a planar two-center instance $S$ is a set of two radius-$r$ disks in the plane whose union covers $S$; a solution for $S$ refers to an $r$-solution with some $r\geq 0$. We denote by $\mathsf{opt}(S)$ the minimum $r$ such that $S$ has an $r$-solution. An $r$-solution $\{D,D^{\prime}\}$ for $S$ is tight if $\mathsf{dist}(D,D^{\prime})$ is minimized (among all $r$-solutions). An $r$-solution $\{D,D^{\prime}\}$ for $S$ is $p$-anchored for a point $p\in\mathbb{R}^{2}$ if $p\in D\cap D^{\prime}$. The anchored two-center problem takes as input a set $S$ of points in the plane and a point $p\in\mathbb{R}^{2}$, and asks for a $p$-anchored $r$-solution for $S$ with the minimum $r$. As discussed in Section 1, Choi and Ahn [8] presented an $O(n\log n)$-time algorithm for the anchored problem.
@@ 3 $\bm{r}$-Coverage
First of all, by the definition of $\alpha_{r}(Q)$, such a disk $D$ must exist for $p\in\partial\alpha_{r}(Q)$. In this proof, we let $q$ be the antipodal point of $p$ on $\partial D$, and show that $q$ lies on $\partial\mathcal{CR}_{r}(Q)$.
@@
Let $e$ be an arc of $\alpha_{r}(Q)$ with two vertices $u$ and $v$. By slightly abusing the notation, define $\tau(e)=\cup_{p\in e}\tau(p)$. Let $D$ be the unique radius-$r$ disk whose boundary contains $e$. Note that $D$ contains $\alpha_{r}(Q)$ and thus contains $Q$ [13, 18, 29]. This property observes that $\tau(e)$ consists of the antipodal points of all points of $e$ on $\partial D$. We call $\tau(e)$ the antipodal arc of $e$. Since we have assumed that the distance of any two points of $Q$ is less than $2r$, $\tau(e)$ does not intersect $e$ [13, 18, 29]. For the reference purpose later, we have the following corollary, following directly from Lemma 6.
@@
We proceed to show that such a disk $D$ is unique. Assume there is another radius-$r$ disk $D^{\prime}$ with $Q\subseteq D^{\prime}$ and $q\in\partial D^{\prime}$. Then the antipodal point $p$ of $q$ on $\partial D$ lies outside of $D^{\prime}$. On the other hand, since $p$ lies on $\partial\alpha_{r}(Q)$ (which has been proved above), $p$ must be contained in $D^{\prime}$ because $D^{\prime}$ contains $Q$ (and thus contains $\alpha_{r}(Q)$). This leads to a contradiction. ∎
@@ 4 Algorithm of Theorem 1
Clearly, $|A|=O(1)$. In other words, $A$ consists of $o$ and all the possible $O(1)$ points of $\mathbb{Z}_{/c}^{2}$ in the radius-$2$ disk $x^{2}+y^{2}\leq{4}$.
@@ 5.1 Geometric lemmas
Let $\{D,D^{\prime}\}$ be a tight $r$-solution for $S$, and $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting the two disk centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. Then, there exist two points $u,v\in(S\cap\partial D)\backslash D^{\prime}$ such that $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$; see Figure 4. Similarly, there exist two points $u^{\prime},v^{\prime}\in(S\cap\partial D^{\prime})\backslash D$ such that $\𝜎\overline{u^{\prime}v^{\prime}}\cap(\ell\backslash\sigma)\neq\emptyset$.
@@
Let $\{D,D^{\prime}\}$ be a set of two radius-$r$ disks, $u,v\in\partial D\setminus D^{\prime}$ be two points whose antipodal points on $\partial D$ are not in $D^{\prime}$, and $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. If $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$, then $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$.
@@
Let $\tau^{\prime}$ be the antipodal arc of $\tau$ on $\partial D$, i.e., $\tau^{\prime}$ consists of the antipodal points on $\partial D$ of all points of $\tau$. Note that $u$ and $v$ are two endpoints of $\tau^{\prime}$. Since $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$, $\tau^{\prime}$ is in a half circle of $\partial D$, and thus $\tau^{\prime}$ must be an arc of the circular hull $\alpha_{r}(\{u,v\})$ of $\{u,v\}$. Applying Corollary 7 with $Q=\{u,v\}$, we obtain that $\tau\subseteq\partial\mathcal{CR}_{r}(\{u,v\})$.
@@ 5.2 Proving the correctness
  • If $r<1$, since $\mathsf{dist}(D,D^{\prime})\leq(2-\frac{5}{c})\cdot r$, there exists a disk $D^{\prime\prime}\subseteq D\cap D^{\prime}$ of radius $1/c$, which in turn contains an axis-parallel square $\Box$ of side-length $1/c$. Note that one of $D$ and $D^{\prime}$ must intersect the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$, because $S\subseteq D\cup D^{\prime}$. As $r<1$, we know that $\Box\subseteq\{(x,y):x^{2}+y^{2}\leq 4\}$. By the same argument as above, $\Box$ contains some point $p\in\{(x,y)\in\mathbb{Z}_{/c}^{2}:x^{2}+y^{2}\leq 4\}\subseteq A$, and thus $\{D,D^{\prime}\}$ is $p$-anchored.